# Resolving the Bulk-Boundary Correspondence Paradox on Low-Symmetry Surfaces of Weyl Semimetals

### v1

@@ Abstract
Low-symmetry surfaces of topological semimetals offer access to boundary phenomena hidden on conventional facets, yet systematic studies remain scarce due to experimental challenges and the lack of a general framework for reconciling bulk and surface periodicities[1, 2, 3, 4, 5]. Here, we investigate the (103) surface of the Weyl semimetal NdAlSi using angle resolved photoemission spectroscopy and density functional theory. The (103) surface is an example of a low symmetry surface that presents an apparent paradox to the bulk-boundary correspondence. The surface periodicity to which the topological surface states are expected to adhere does not correspond to the bulk periodicity of the Weyl points. By showing that successive bulk Brillouin zones generate replicas that accumulate into a superlattice commensurate with the surface Brillouin zone, we demonstrate how the apparent bulk boundary correspondence paradox is resolved and establish a universal criterion for arbitrary facets. The framework and experimental results further suggests that overlapping Fermi arcs can hybridize into closed Fermi arc loops, enriching boundary topology and enabling unconventional transport, interference, and collective phenomena unique to low-symmetry facets.
@@ Introduction
Topological semimetals have emerged as a fertile ground for exploring exotic quantum states of matter, where nontrivial band topology gives rise to robust surface excitations and novel bulk transport phenomena[1, 2, 3]. Among them, Weyl semimetals are characterized by linear bulk band crossings–Weyl points–that act as monopoles of Berry curvature in momentum space[1, 2, 3, 6, 7, 8, 9, 10, 11]. A defining feature of Weyl semimetals is the bulk-boundary correspondence (BBC), which dictates that pairs of Weyl points of opposite chirality in the bulk are connected by open Fermi arc states on the surface[4, 5]. This correspondence has been extensively established on high-symmetry surfaces such as (001) and (100), where the surface Brillouin zone (BZ) is commensurate with the bulk projection[7, 8, 9, 12, 13, 14, 16, 15].
In contrast, low-symmetry surfaces remain largely unexplored, even though they are uniquely positioned to reveal fundamentally new physics. Their intrinsic incommensurability between surface and bulk BZ periodicities raises a conceptual puzzle: while Weyl points project according to the bulk BZ, surface Fermi arcs follow the surface BZ (SBZ), this mismatch suggesting an apparent breakdown of the BBC. At the same time, low-symmetry facets naturally disentangle surface states and surface-bulk projected states (SBPSs) from real bulk states. On high-symmetry surfaces, rotational and mirror symmetries enforce strong constraints and cancellations that suppress many boundary responses. By contrast, on low-symmetry surfaces these constraints are lifted: Fermi arcs can become highly anisotropic or undergo reconstruction[14, 17, 18], parallel arc segments can enhance nesting-driven instabilities[19, 20, 21, 22], Berry-curvature contributions no longer cancel and instead give rise to nonlinear and nonreciprocal responses[23, 24, 25, 26], and the reduction of crystalline symmetries facilitates the emergence of higher-order boundary states[27]. Despite this promise, experimental and theoretical studies of low-symmetry surfaces have been scarce, owing to the experimental difficulty of preparing and interrogating non-natural cleavage planes and the lack of a general theoretical framework for reconciling bulk and surface periodicities on such facets[1, 2, 3, 4, 5].
Here we address this challenge by investigating the low-symmetry (103) surface of the Weyl semimetal NdAlSi using angle resolved photoemission spectroscopy (ARPES) and density functional theory (DFT). By combining disorder-induced suppression of coherent surface states on low-symmetry (103) facets, we introduce a new experimental strategy that enables the direct and unambiguous identification of surface-bulk resonance states (SBRSs)[28, 29, 30, 31, 32, 33, 34, 35] in a Weyl semimetal. At the same time, we demonstrate that the apparent paradox of BBC is resolved when bulk-to-surface projections are extended beyond the first BZ: higher-order projections generate translated replicas whose superposition yields a superlattice periodicity commensurate with the SBZ. This establishes a generalized framework for BBC on arbitrary crystal facets, demonstrating that the apparent breakdown on low-symmetry surfaces is not a failure of the BBC but a consequence of incomplete bulk projection. The resulting least common multiple (LCM) criterion provides a universal means of reconciling bulk and surface states and opens a pathway for engineering exotic boundary phenomena across a broad class of topological materials. More broadly, our results open a new path in the study of low-symmetry surfaces of topological semimetals, laying the groundwork for systematic exploration of their hidden boundary physics. By showing how such facets naturally disentangle and amplify the interplay between surface and bulk states, our findings create opportunities to design unconventional responses and collective excitations that remain inaccessible on conventional high-symmetry surfaces.
@@ Results
To address above issues, we turn to the Weyl semimetal NdAlSi[13, 14, 15, 22], a member of the non-centrosymmetric RAlSi (R=rare earth) family[36, 37, 38, 39] that provides an excellent platform for studying low-symmetry facets. In particular, we investigate the (103) surface, which is never explored yet offers a unique opportunity to directly test the validity of BBC. The crystallographic structure of NdAlSi and the orientation of the (103) surface are shown in Fig. 1a. This cut exposes a low-symmetry plane that is rarely studied in ARPES due to the experimental challenges associated with surface preparation. Fig. 1b illustrates the construction of the bulk BZ and its projection onto the (103) surface. Unlike high-symmetry surfaces such as (001) or (100), the (103) surface projection produces a fundamental mismatch: the periodicity of the SBZ (green lines) does not coincide with that of the projected bulk BZ (pink lines), as shown in Fig. 1c.
The surface orientation was verified by X-ray diffraction (XRD), which confirms the high-quality exposure of the (103) plane (Fig. 1d) (see Methods for sample preparation). Fig. 1e and 1f display Fermi surfaces of NdAlSi measured at 150 eV on well defined (Fig. 1e) and disordered (Fig. 1f) regions of the (103) surface. On the well defined area, the measured Fermi surface (Fig. 1e) is dominated by surface states (see Section 1 in Supplementary Materials for details). On the disordered region, however, the surface states are almost suppressed[15], revealing additional features that cannot be assigned to pure surface bands. Fig. 1e clearly shows that the surface states respect the periodicity of the SBZ. In contrast, Fig. 1f reveals that the bulk states exhibit a periodicity that does not coincide with the SBZ. This provides direct experimental evidence of the intrinsic non-commensurability between the surface states and bulk state projections on the (103) surface. While not unexpected for topologically trivial systems, this observation raises a central question for Weyl semimetals: If Weyl points follow the bulk BZ periodicity while surface Fermi arcs follow the SBZ periodicity, does the BBC break down on low-symmetry surfaces?
To understand the nature of the electronic states on the (103) surface and answer the above question, we performed systematic ARPES measurements combined with surface-projected DFT calculations based on Green’s function methods. Fig. 2a shows the Fermi surface measured at 41 eV photon energy on a well defined region of the surface, which is dominated by surface states (see Section 2 in Supplementary Materials for results under different polarizations). The corresponding DFT calculation for a Si-terminated surface (Fig. 2b) reproduces the observed features quite well, confirming their surface origin. Data collected on a disordered region of the surface under the same conditions show that the surface states are strongly suppressed (Fig. 2c)[15]. This suppression acts as a natural “filter”, allowing otherwise hidden SBPSs to emerge. When a surface state enters the bulk-projected region, it hybridizes with the bulk continuum and evolves into a SBRS[28, 29, 30, 31, 32, 33, 34, 35], resulting in hybridized features observable in ARPES (see Section 3 in the Supplementary Materials). The DFT calculation reproduces these features, as shown in Fig. 2b. Consistently, Green’s function calculations capture both sharp surface states and additional bulk-projected features.
Momentum-resolved band dispersions along two representative cuts (Cut1 and Cut2) are shown in Figs. 2e and 2f. These sharp dispersive bands agree well with surface-projected DFT calculations (Figs. 2g and 2h). By contrast, along the same cuts on the disordered surface (Figs. 2i and 2j), broader and bulk-like bands appear, consistent with calculated SBPSs (Figs. 2k and 2l). Sections 3 and 4 of the Supplementary Materials provides more details on the identification of surface states, bulk states, SBPSs and SBRSs.
@@
Having established the surface states and SBPSs, we next turn to the bulk electronic structure underlying the (103) surface. The 3D bulk Fermi surface from DFT is shown in Fig. 3a, while Fig. 3b illustrates the (103) surface orientation and its perpendicular cross-section in momentum space. Calculated bulk Fermi surface slices along the [103] direction (k⟂) and perpendicular to the (103) surface (cyan plane in Fig. 3b) are presented in Figs. 3c and 3d, serving as theoretical benchmarks.
@@
Experimentally, the photon energy dependent ARPES spectral intensity map at the Fermi level along ky (Fig. 3e) reveals strong k⟂ modulation, a hallmark of bulk-derived states. Systematic Fermi surface measurements over photon energies from 125 eV to 195 eV (Figs. 3f-i, 3n-q) further confirm this k⟂ dispersion. The corresponding DFT calculations (Figs. 3j-m, 3r-u) reproduce the experimental features with remarkable accuracy. More detailed bulk Fermi surfaces of the NdAlSi (103) surface are provided in Section 5 of the Supplementary Materials.
@@
To clarify how this resolves the apparent BCC paradox, we track how the bulk-projection periodicity evolves as higher-order BZs are included. If only the first bulk BZ is considered, the projected Weyl points (Fig. 4a) clearly mismatch the theoretical SBZ (purple lines) and the experimental SBZ (green lines). Adding the second-order (Fig. 4b) and third-order BZs and Weyl points (Fig. 4c) along progressively restores the periodicity, bringing the bulk projection closer to the surface states. This restoration follows a LCM criterion: the projected displacement between successive bulk BZs closes only after an integer multiple of the in-plane reciprocal vectors. For the (103) surface of NdAlSi, the minimum closure occurs after steps (see Section 6 in the Supplementary Materials). As illustrated in Fig. 4d, the projected Weyl points recover the theoretical SBZ periodicity (purple lines) once higher-order BZs are included.
@@
Beyond this restoration, the higher‐order projections reveal an emergent smaller repeating unit: a “new SBZ” (green lines in Fig. 4d) whose period along is one third of the theoretical SBZ. This “new SBZ” coincides with the one observed experimentally. Its origin lies in the fact that the bulk BZs of three consecutive, nonequivalent (103) surface unit cells differ by a phase of 2/3 in momentum space; their hybridization produces the “new smaller SBZ”. Equivalently, this effect can be viewed as a fixed lateral shift between successive bulk projections (see Section 6 in the Supplementary Materials), which manifests as a momentum-space moir pattern–a weak superlattice modulation with a period one third of the original SBZ. In general, whether a single bulk BZ suffices depends on surface symmetry. For high-symmetry surfaces such as (001), the first bulk BZ already coincides with the SBZ (Fig. 4f). For low-symmetry surfaces such as (103), however, successive bulk projections are laterally shifted (Fig. 4g), and only after steps does the bulk periodicity close. Exact spectral reconstruction is achieved only in the infinite-projection limit (see Section 6 in the Supplementary Materials). Thus, the observed mismatch is not a breakdown of BBC but rather the natural consequence of incomplete bulk projection, while the emergent new SBZ reflects the moir-like interference among multiple shifted projections.
This insight also clarifies why different computational schemes perform so differently. Finite-slab calculations (six layers in our case) truncate the bulk projections and therefore miss the SBPSs that are essential on low-symmetry facets. By contrast, semi-infinite surface Green’s function calculations include contributions from infinitely many bulk BZs and thereby capture both surface states and SBPSs in excellent agreement with experiment. Thus, while slab models may suffice for commensurate high-symmetry surfaces[13, 14], Green’s function methods are indispensable for correctly describing the electronic structure of low-symmetry surfaces such as (103) (see Section 4 in the Supplementary Materials).
@@
An additional subtlety arises when considering the surface Fermi arcs themselves. On high-symmetry surfaces such as (001), arcs from successive bulk BZs project in perfect registry, perhaps typically preserving the canonical picture of open arcs that directly connect projected Weyl points (Fig. 4f). In contrast, on low-symmetry surfaces like (103), the incommensurate periodicities of bulk and SBZs cause consecutive arcs to project with mismatched positions. This misalignment naturally drives hybridization between arcs originating from different bulk zones (Fig. 4g). Consequently, the observed arcs are not simple replicas but collective superpositions, and in certain cases multiple arcs can merge into closed surface Fermi arc loops (SFALs)–a boundary topology inaccessible on high-symmetry facets (Fig. 4d, see Section 7 of the Supplementary Materials for details). While our framework and present ARPES superposition features suggest the emergence of SFALs, their unambiguous demonstration remains open. Future high-resolution spin ARPES or quantum oscillation studies will be key to confirming their presence. Such loop-like structures enrich the boundary phenomenology of Weyl semimetals: their closed topology provides natural -space trajectories for novel oscillatory and interference effects, while their overlapping geometry may promote nesting-driven instabilities or correlated orders.
By investigating the low-symmetry (103) surface of the Weyl semimetal NdAlSi, we introduce a new experimental route to directly identify SBRSs. This is enabled by disorder-induced suppression of surface states, which acts as a natural filter to distinguish the SBRS from others. Furthermore, the (103) surface highlights an apparent paradox of the BBC. Contrary to the topological expectation that Fermi arcs and Weyl points should share the same periodicity, the Weyl points follow the bulk while the Fermi arcs display the surface periodicity. We resolve this puzzle by showing that the mismatch originates not from a break down of the BBC but from incomplete bulk projection. Successive BZ Weyl points generate translated surface Fermi arcs whose superposition yields a superlattice commensurate with the surface, thereby restoring the BBC. This LCM construction can be viewed as a momentum-space moir modulation, the k-space analogue of real-space moir stripes, where long-period modulations govern the reconstruction of boundary states.
More broadly, our results inaugurate a new frontier in the study of low-symmetry surfaces. The possibility of SFALs enriches the boundary topology of Weyl semimetals, suggesting unconventional transport, interference, and collective phenomena beyond the canonical open-arc paradigm. The conceptual framework extends beyond Weyl semimetals to Dirac and nodal-line semimetals, higher-order topological systems, and even non-centrosymmetric superconductors. Its analogy to moir physics in stacked two-dimensional materials further highlights its generality: in both real space and momentum space, mismatched periodicities generate emergent superlattices that control hybridization and boundary states. The same principle also bridges to quasicrystals, where intrinsic incommensurability requires infinite projection to restore quasi-periodicity. A more detailed discussion is provided in Section 8 of the Supplementary Materials. In summary, our work resolves the apparent BBC paradox on low-symmetry surfaces and provides a unifying momentum-space moir framework connecting crystalline and quasicrystalline boundaries, opening a platform to uncover hidden boundary phenomena and engineer exotic responses in quantum matter.

### v3

# Hidden Moiré Topology of Low-Symmetry Weyl Surfaces

@@ Abstract
Topological materials are defined by the correspondence between bulk topology and boundary states, yet this correspondence becomes enigmatic on low-symmetry surfaces where bulk and surface periodicities are inherently mismatched. Here we reveal a hidden moiré topology emerging on the (103) surface of the Weyl semimetal NdAlSi. Angle-resolved photoemission spectroscopy uncovers closed Fermi-arc loops and momentum-space moiré modulations, phenomena unanticipated in conventional topological theory. We show that these emerge from incomplete bulk projection and multi-cell interference governed by a least-common-multiple framework. Least-common-multiple guided DFT and Green’s-function calculations quantitatively reproduce the observed spectra, establishing the universality of this commensuration rule. These findings transform a long-standing paradox of bulk-boundary correspondence into a new paradigm of momentum-space moiré reconstruction, bridging crystalline and quasicrystalline topologies and opening routes to flat-band engineering on complex surfaces.
@@ Introduction
Since the discovery of topological materials, a central question in condensed-matter physics has been how bulk topology manifests at material boundaries. Nowhere is this connection more striking than in Weyl semimetals[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], where open Fermi arcs on high-symmetry surfaces such as (001)/(100) of tetragonal or orthorhombic lattices[5, 6, 7, 12, 13, 14] and the (0001) facet of hexagonal lattices[15, 16] directly connect the surface projections of bulk Weyl points. These canonical surfaces, where bulk and surface periodicities are trivially commensurate, have accordingly dominated both experimental and theoretical studies of topological semimetals.
In contrast, low-symmetry surfaces remain largely unexplored[17, 18, 19], despite offering a qualitatively new regime of topology. Their intrinsic incommensurability between bulk and surface Brillouin zones (SBZs) disrupts the conventional one-to-one mapping between bulk Weyl points and surface Fermi arcs, raising the question of how bulk-boundary correspondence (BBC) manifests when translational commensurability is lost. Whether BBC in this regime survives in a modified form or requires a new organizing principle remains unknown, as no direct experimental observation or coherent theoretical framework has yet addressed it. This gap stems from formidable challenges on both fronts: experimentally, low-symmetry facets rarely cleave into atomically uniform planes[19], and even when achieved, disorder and multiple terminations obscure intrinsic surface states in angle-resolved photoemission spectroscopy (ARPES) measurements; theoretically, conventional slab models assume commensurate periodicities and therefore fail to capture the correct surface repeat unit, making density functional theory (DFT) reproduction of experimental spectra notoriously difficult. Because Weyl semimetals themselves originate from broken symmetries, low-symmetry surfaces naturally extend this principle into the boundary, where reduced crystalline constraints are expected to generate qualitatively new boundary phenomena–ranging from anisotropic or reconstructed Fermi arcs[13, 20, 21] to nonlinear transport[26, 27, 28, 29] and higher-order boundary states[30]. A framework capable of reconciling bulk-surface incommensurability–both conceptually and computationally–is therefore essential for advancing this frontier.
Here, we overcome these barriers by realizing a well-defined (103) surface of the Weyl semimetal NdAlSi through precise pre-orientation and pre-cutting of single crystals, followed by controlled cleavage that exposes atomically uniform terraces. This preparation enables high-resolution ARPES measurements that reproducibly resolve multiple surface-state configurations. The spectra reveal an apparent mismatch between bulk-projected Weyl points and the SBZ–a direct manifestation of bulk-surface incommensurability. Guided by this observation, we show that the mismatch is not a breakdown of BBC but a natural outcome of incomplete bulk projection. By extending bulk-to-surface mapping beyond the first BZ, we establish a least common multiple (LCM) criterion that restores bulk-surface commensuration and explains the emergence of a “new smaller SBZ”, corresponding to the momentum-space moiré pattern observed experimentally. For NdAlSi (103), surface-projected DFT calculations based on this framework quantitatively reproduce the ARPES spectra and capture the characteristic interference among inequivalent surface cells. These results demonstrate that the observed moiré modulation is the direct consequence of LCM-guided reconstruction, where finite phase offsets between neighboring surface cells generate long-period interference and hybridization between surface and bulk-projected states. More broadly, this universal framework resolves the apparent BBC paradox and establishes low-symmetry surfaces as momentum-space analogues of moiré systems, linking incommensurability, surface reconstruction, and correlated boundary phenomena across a broad class of quantum materials.
@@ Results
We begin with the Weyl semimetal NdAlSi[12, 13, 14, 25], a non-centrosymmetric member of the RAlSi (R = rare-earth) family[31, 32, 33, 34], which hosts well-characterized Weyl points and high-quality single crystals ideally suited for surface-sensitive spectroscopy. We focus on the low-symmetry (103) surface, which has never been explored yet provides a natural setting to probe how BBC manifests when bulk and surface periodicities are incompatible. The crystallographic structure and orientation of the (103) surface are shown in Fig. 1a, while Fig. 1b illustrates the bulk BZ and its projection. Unlike high-symmetry facets such as (001) or (100), the (103) projection exhibits an intrinsic mismatch between the periodicities of the SBZ (purple lines) and the projected bulk BZ (pink lines), as shown in Fig. 1c–this incommensurability forms the experimental starting point of our study.
The intrinsic incommensurability revealed by geometric projection motivates direct experimental verification on an atomically defined (103) surface. The surface orientation was confirmed by X-ray diffraction (XRD), which verifies the high-quality exposure of the (103) plane (Fig. 1d; see Methods for sample preparation). Fermi-surface maps acquired at 150 eV further highlight the crucial role of surface quality: uneven regions (Fig. 1e) show diffuse features dominated by bulk states, whereas well-defined terraces (Fig. 1f) exhibit sharp surface states that follow the periodicity of the surface Brillouin zone (SBZ, purple lines). A closer inspection, however, reveals an additional modulation with a smaller period (green lines), indicating a more subtle reconstruction. On the uneven region, the suppression of coherent surface states[14] exposes bulk-like bands whose periodicity does not coincide with that of the SBZ. These observations demonstrate that the mismatch between bulk and surface periodicities is not merely a geometric construct but a directly observable phenomenon on the (103) surface. This raises a central question: if Weyl points follow the bulk BZ periodicity while surface Fermi arcs adhere to the SBZ, does the BBC break down on low-symmetry surfaces?
@@
To understand the nature of the electronic states on the (103) surface and answer the above question, we performed systematic ARPES measurements combined with bulk DFT calculations. The 3D bulk Fermi surface from DFT is shown in Fig. 2a, while Fig. 2b illustrates the (103) surface orientation and its perpendicular cross-section in momentum space. Calculated bulk Fermi surface slices along the [103] direction (k⟂) and perpendicular to the (103) surface (cyan plane in Fig. 2b) are presented in Figs. 2c and 2d, serving as theoretical benchmarks.
@@
Experimentally, the photon energy dependent ARPES spectral intensity map at the Fermi level along ky (Fig. 2e) reveals strong k⟂ modulation, a hallmark of bulk-derived states. Systematic Fermi surface measurements over photon energies from 125 eV to 195 eV (Figs. 2f-i, 2n-q) further confirm this k⟂ dispersion. The corresponding DFT calculations (Figs. 2j-m, 2r-u) reproduce the experimental features with remarkable accuracy. More detailed bulk Fermi surfaces of the NdAlSi (103) surface are provided in Section 3 of the Supplementary Materials.
@@
To clarify how this resolves the apparent BCC paradox, we track how the bulk-projection periodicity evolves as higher-order BZs are included. If only the first bulk BZ is considered, the projected Weyl points (Fig. 3a) clearly mismatch the theoretical SBZ (purple lines) and the experimental SBZ (green lines). Adding the second-order (Fig. 3b) and third-order (Fig. 3c) BZs and Weyl points along progressively restores the periodicity, bringing the bulk projection closer to the surface states. This restoration follows a LCM criterion: the projected displacement between successive bulk BZs closes only after an integer multiple of the in-plane reciprocal vectors. For the (103) surface of NdAlSi, the minimum closure occurs after steps (see Section 4 in the Supplementary Materials). As illustrated in Fig. 3d, the projected Weyl points recover the theoretical SBZ periodicity (purple lines) once higher-order BZs are included.
@@
Beyond this restoration, the higher‐order projections reveal an emergent smaller repeating unit: a “new SBZ” (green lines in Fig. 3d) whose period along is one third of the theoretical SBZ. This “new SBZ” coincides with the one observed experimentally. Equivalently, this effect can be viewed as a fixed lateral shift between successive bulk projections (see Section 4 in the Supplementary Materials), which manifests as a momentum-space moiré pattern–a weak superlattice modulation with a period one third of the original SBZ (Fig. 3d). In general, whether a single bulk BZ suffices depends on surface symmetry. For high-symmetry surfaces such as (001) of NdAlSi, the first bulk BZ already coincides with the SBZ (Fig. 3f). For low-symmetry surfaces such as (103), however, successive bulk projections are laterally shifted (Fig. 3g), and only after steps does the bulk periodicity close. Exact spectral reconstruction is achieved only in the infinite-projection limit (see Section 4 in the Supplementary Materials). Thus, the observed mismatch is not a breakdown of the BBC but rather the natural consequence of incomplete bulk projection, while the emergent new SBZ reflects the moiré-like interference among multiple shifted projections. Here we display one representative surface-state configuration; additional terminations observed on the same cleaved surface are presented in Section 5 of the Supplementary Materials.
@@
An additional subtlety arises when considering the surface Fermi arcs themselves. On high-symmetry surfaces such as (001) of NdAlSi, arcs from successive bulk BZs project in perfect registry, typically preserving the canonical picture of open arcs directly connecting projected Weyl points (Fig. 3f). In sharp contrast, on low-symmetry surfaces like (103), the incommensurate periodicities of bulk and SBZs cause consecutive arcs to project with lateral mismatches. This misalignment naturally drives hybridization between surface arcs originating from bulk zones that are displaced both in-plane and out-of-plane (Fig. 3g), so that the observed features are not simple replicas but collective superpositions. In certain cases, multiple arcs merge into closed SFALs–a boundary topology unusual on high-symmetry facets (Fig. 3d; see Section 6 of the Supplementary Materials for details).
Our framework, together with the best ARPES resolution presently achievable, provides compelling evidence for the emergence of SFALs. These loop-like states enrich the boundary phenomenology of Weyl semimetals: their closed -space trajectories provide natural channels for quantum interference and oscillatory responses, while the concentrated Berry curvature they enclose may enhance nonlinear or nonreciprocal transport. Moreover, overlapping arc segments can promote nesting instabilities, offering a route to correlated boundary orders. Future high-resolution probes such as spin-resolved ARPES or quantum oscillations will provide complementary confirmation and deeper insight into the microscopic structure of these loops.
To capture such arc reconstruction and hybridization theoretically, the surface periodicity used in modeling must be commensurate with bulk projections. This requirement is fulfilled by the LCM criterion, which provides a systematic prescription for selecting appropriate surface unit cells in DFT slab calculations to ensure commensurability between bulk projections and surface periodicities. In the case of the (103) surface of NdAlSi, the minimal commensuration requires , corresponding to a three-cell surface supercell composed of three consecutive unit cells. By contrast, a single-cell slab (or restricting to the first bulk BZ) systematically fails to reproduce the experimental periodicity, because the projection closure remains incomplete. With the correct three-cell supercell, surface-projected DFT captures the reduced periodicity and dominant features of the ARPES spectra. For a Si-terminated (103) surface (Fig. 4b), the LCM-consistent calculation reproduces the emergent “new smaller SBZ”, originating from the phase offsets of between bulk BZs of three consecutive, nonequivalent (103) surface unit cells. Their hybridization yields the reduced periodicity observed in experiment.
Despite this overall agreement, finer aspects of the Fermi-arc connectivity are not fully reproduced in the idealized calculations. In ARPES, closed surface Fermi-arc loops (SFALs) appear within the reduced SBZ, whereas the LCM-consistent DFT yields only open arcs. This discrepancy most likely reflects two limitations of the model: first, the absence of zigzag-type reconstructions or mixed local terminations, which in real samples enhance coupling between surface and bulk states; and second, the incomplete treatment of hybridization across multiple projected bulk BZs, whose lateral mismatch further promotes loop-like connectivity. We therefore regard the observed SFAL as an emergent property of low-symmetry surfaces beyond what is captured by ideal slab calculations. In addition to the Si-terminated regions, ARPES measurements also reveal two further surface-state configurations on other areas of the same cleaved (103) surface (see Fig. S10 in the Supplementary Materials). These distinct spectra share the reduced periodicity but do not match any of the terminations tested in Fig. S11 of the Supplementary Materials, suggesting a more complex origin such as zigzag-type reconstructions or mixed local terminations beyond the simple three-cell picture in Fig. S9b in the Supplementary Materials. A detailed discussion of these additional states, and their relation to the LCM criterion, is provided in Section 5 of the Supplementary Materials.
Beyond the ordered regions, disorder offers a complementary perspective on the same framework. When spectra are collected on a disordered area of the (103) surface, sharp surface states are strongly suppressed (Fig. 4c)[14], effectively acting as a natural “filter” that allows otherwise hidden surface bulk-projected states (SBPSs) to emerge. Remarkably, these SBPSs are well reproduced by the same surface-projected DFT calculations (Fig. 4b), underscoring the robustness of the LCM-guided modeling. When a surface state enters the bulk-projected region, it hybridizes with the bulk continuum and evolves into a surface bulk-resonant state (SBRS)[35, 36, 37, 38, 39, 40, 41, 42], producing hybridized features observable in ARPES. Momentum-resolved band dispersions along two representative cuts (Cut1 and Cut2) are shown in Figs. 4e and 4f. These sharp dispersive bands agree well with surface-projected DFT calculations (Figs. 4g and 4h). By contrast, along the same cuts on the disordered surface (Figs. 4i and 4j), broader and bulk-like bands appear, consistent with calculated SBPSs (Figs. 4k and 4l). Section 8 of the Supplementary Materials provide more details on the identification of surface states, bulk states, SBPSs, and SBRSs.
Discussion and conclusion
By investigating the low-symmetry (103) surface of the Weyl semimetal NdAlSi, we uncover a previously hidden regime of topological boundary reconstruction that fundamentally extends the conventional notion of BBC. While Weyl points follow the intrinsic bulk periodicity, the surface Fermi arcs adhere to a reduced periodicity defined by the SBZ. What initially appears as a paradox of the BBC is in fact a consequence of incomplete bulk projection, where successive bulk BZs generate laterally phase-shifted replicas that close only after multiple unit cells. This reconstruction restores commensuration through a LCM criterion, naturally producing the emergent “new smaller SBZ” observed in ARPES. The resulting interference pattern manifests as a momentum-space moiré modulation, in which neighboring surface unit cells contribute discrete phase offsets analogous to twisted or displaced layers in real-space moiré systems.
This momentum-space moiré analogy carries profound implications. The periodic phase interference between inequivalent surface unit cells can compress surface dispersions and generate flat-band-like electronic structures, thereby amplifying correlation effects at the boundary. In this picture, low-symmetry surfaces emerge as natural moiré analogues in momentum space, where the interplay of topology, symmetry, and interference gives rise to boundary states beyond the canonical open-arc paradigm. The observation of surface Fermi-arc loops (SFALs) further demonstrates that low-symmetry surfaces host boundary topologies inaccessible on conventional facets, pointing to a new class of topological states governed by multi-cell commensuration and hybridization.
More broadly, the LCM framework provides a unified prescription for reconciling bulk and surface periodicities across complex crystals. Its mathematical structure generalizes naturally to Dirac and nodal-line systems, higher-order topological phases, and non-centrosymmetric superconductors. Beyond crystalline materials, the same principle extends to quasicrystals and aperiodic systems, where commensuration becomes infinite and quasi-periodicity replaces translational order. From a practical standpoint, the LCM criterion offers a robust guideline for determining the minimal surface supercell required for accurate DFT slab calculations–long a bottleneck in modeling low-symmetry surfaces.
In summary, our work transforms a perceived inconsistency of the BBC into a new paradigm of boundary reconstruction, establishing a bridge between topological surface physics and moiré interference phenomena. By revealing that incommensurability itself can serve as a design principle for emergent boundary states, these findings open pathways toward flat-band formation, enhanced correlations, and topological interference effects in both crystalline and quasiperiodic systems.